OpenStudy (anonymous):

Let g(x) be the inverse of f(x)=x^{3}+3x+5. Calculate g(9) [without finding a formula for g(x)] and then calculate g'(9). g(9) = 1 g'(9) = ?

myininaya (myininaya):

f(x)=y implies g(y)=x since g and f are inverses so we need to find when f(x)=9

myininaya (myininaya):

That is you need to solve x^3+3x+5=9 and you technically don't have to solve just trial and error it

myininaya (myininaya):

usually these are made so the guessing is easy

myininaya (myininaya):

and it is pretty easy in this case

OpenStudy (anonymous):

1 right?

myininaya (myininaya):

1 is right 1+3+5=9

myininaya (myininaya):

so f(1)=9 so f^(-1)(9)=1 (or what they called it g(9)=1)

myininaya (myininaya):

\[\frac{d}{dx} f^{-1}(x)=\frac{1}{f'(f^{-1}(x))}\]

myininaya (myininaya):

\[\frac{d}{dx}g(x)|_{x=9}=\frac{1}{f'(g(9))}\]

myininaya (myininaya):

So now find f'

myininaya (myininaya):

and plug 1 into it

OpenStudy (anonymous):

so \[f'=3x^2+3. \] plug in 1 you get 6 then plug that into 1/f'(g(9)) so the answer is 1/6. I think.

myininaya (myininaya):

Sounds find.